find the length of the shortest side of a triangle whose perimeter is 66 and the sides are marked as x+5, 3X-1 AND 4X-2 what is the shortest side
Found 2 solutions by macston, ikleyn:
Answer by macston(5194)

.
Perimeter is the sum of the sides.
.
(x+5)+(3x-1)+(4x-2)=66
8x+2=66
8x=64
x=8
.
x+5=8+5=13
The x+5 side is 13
.
3x-1=3(8)-1=24-1=23
The 3x-1 side is 23
.
4x-2=4(8)-2=32-2=30
The 4x-2 side is 30
.
ANSWER: The shortest side is x+5=13
.

Answer by ikleyn(52800)

.
find the length of the shortest side of a triangle whose perimeter is 66 and the sides are marked as x+5, 3X-1 AND 4X-2 what is the shortest side
---------------------------------------------------------------------
Your equation is 

(x+5) + (3x-1) + (4x-2) = 66.

8x + (5-1-2) = 66,

8x + 2 = 66,

8x = 66 - 2 = 64,

x = 8.

The sides of the triangle are:

side 1:  x + 5 = 8+5      = 13;
side 2:  3x-1  = 3*8-1    = 23;
side 3:  4x - 2 = 4*8 - 2 = 30,

The shortest side is 13.